JÓZSEF ATTILA UNIVERSITY  Aradi vértanúk tere 1.

Szeged, Hungary–6720

BOLYAI INSTITUTE  Tel/Fax: (36) 62 326246

Conference on Lattices and Universal Algebra

Szeged, Hungary (August 3–7, 1998)

FIRST ANNOUNCEMENT

Dear Colleagues,

We cordially invite you to the nineteenth algebraic conference held in Szeged. The conference will start on August 3 (Monday) morning, and will conclude on August 7 (Friday) early afternoon. The participants are expected to arrive on August 2.

The scientific program will consist of ten 50-minute plenary lectures on current research in lattice theory and universal algebra, a number of 20-minute contributed talks in parallel sections, and a problem section. The following professors have already accepted our invitation to give a plenary lecture: B. Davey (Bundoora, Australia), C. Herrmann (Darmstadt, Germany), E. W. Kiss (Budapest, Hungary), V. Gorbunov (Novosibirsk, Russia), G. Grätzer (Winnipeg, Canada), M. Valeriote (Hamilton, Canada), F. Wehrung (Caen, France) and R. Willard (Waterloo, Canada). We expect about 90 participants.

Registration form. If you are interested in our conference, please fill in and return the enclosed registration form by May 15, preferably via e-mail. (Before e-mailing, please follow the instruction given in the first ten lines of the file.)

Abstracts. Participants who intend to give a talk are kindly asked to send an abstract by May 31. The abstract should include the author’s name and affiliation below the title. The length of the abstract should not exceed one page. The page size should be 21.5 cm × 28 cm (8.5” × 11”) with 2.5 cm (1”) margins on all four sides. We accept abstracts sent by e-mail if written in Plain TEX. In case of other variants of TEX the dvi file is also requested. When figures occur or the document was not written in TEX, we will hopefully be able to handle postscript files. When you submit your abstract via e-mail, please send the file separately from other correspondence. We also accept camera-ready typed or printed abstracts. Abstracts received by deadline May 31 will be collected in a booklet. Copies of the booklet will be distributed among participants upon arrival.

Proceedings. Algebra Universalis will dedicate a separate issue for the best papers in connection with the conference lectures. The papers will be refereed according to Algebra Universalis standards. Papers can be submitted during the conference. The deadline of the submission is September 30.

Social events. A banquet will be held on Monday evening. The problem session will be combined with a small comfortable excursion. All registered participants and accompanying persons are entitled to participate in the banquet and the excursion.

Registration. For participants the registration fee is US $ 50 before May 15 and US $ 70 thereafter. For accompanying persons the registration fee is US $ 30. Children under 6 pay no fee.

Accommodation. Reservations are handled by the Travel Agency ”Congress & Hobby Service”, cf. the Registration and Order Form for details. (Your Registration and Order Form will be passed to this agency.) You can book a room by indicating your choice on this form and sending a \underline{deposit} of US $ 20, US $ 40 or US $ 100, as indicated. Reservations without deposits will not be accepted.

Payment. The registration fee and ho(s)tel deposit are to be paid by money order/certified cheque made payable to

József Attila University (LatticeConf/Czédli)

or by a bank transfer to

name of bank: Postabank és Takarékpénztár Rt. address of bank: Szeged, Hungary-6720

account no.: JATE: 026–716342–USD–941–000

to: LatticeConf/Czedli.

The name(s) of sender(s) and LatticeConf/Czedli should not be omitted! In exceptional cases the registration fee might be paid upon arrival. But then personal cheques will not be accepted and change will be available only in Hungarian forints. If you cancel your participation before June 30, the whole amount minus bank costs will be refunded. There is no refund if you cancel it after July 1.

Travel information. The conference will take place in the building of the Bolyai Institute, downtown in Szeged. Szeged is in the southeastern part of Hungary, approx. 200 kilometers from Budapest. The train connection between Szeged and Budapest is good. Trains leave approximately in every two hours in daylight. Up-to-date travel information will be given in the second announcement. Note that in order to catch the last train to Szeged, you should exit from Budapest Airport by 6 p.m. or so, provided you arrive in Hungary by plane.

Deadlines. The Registration and Order Form and the ho(s)tel deposits have to be here by May 15. May 15 is also the deadline for the registration fee. Abstracts should arrive by May 31. Refund is possible only in case of cancellation not later than June 30. Papers can be submitted by September 30.

Miscellaneous. There will be no organized meals. There are restaurants nearby where a lunch or dinner costs about US $ 3 – 7. A lunch in the university restaurant is about US $ 2. We are applying for some local funds. Depending on the success of our applications, financial support — ranging from reducing the registration fee to a contribution to living or travel costs — might be provided upon request for some participants. It depends on the participant’s citizenship whether he/she needs a visa. A visa can be obtained at any Hungarian Embassy or Consulate. The second announcement of the conference containing detailed travel information will be sent by July 15. We are indebted to Algebra Universalis for the mailing list that we have built into ours.

Szeged, March 15, 1998.

G. Czédli  E. K. Horváth  G. Takách  L. Zádori

(the organizing committee)

Conference on Lattices  REGISTRATION AND ORDER FORM
and Universal Algebra  Return by May 15, 1998 to algebra@math.u-szeged.hu
Szeged, August 3–7, 1998  or mail to G. Czédli (see address above)

Surname: First name(s): , ◯ male / ◯ female Affiliation: Mailing address:

Email: Phone/Fax: Accompanying person(s): + children under 6 Title of talk:

Abstract  ◯ is enclosed  ◯ will be sent by May 15.

I order the following services from CONGRESS & HOBBY SERVICE:

ACCOMMODATION

for  person(s), for  nights
from August  (arrival) to August  (departure)  Daily room rate for
1 person 2 persons 3 persons
Student’s Hostel “Eötvös” (deposit: US $20)  ◯ US $10 ◯ US $12  —

100 meters from conference site, rooms with running water

Student’s Hostel “Apáthy” (deposit: US $20)  ◯ US $15 ◯ US $20 ◯ US $27

300 meters from conference site, rooms with private bathroom

Hotel TISZA∗∗ (deposit: US $40)  ◯ US $35 ◯ US $40 ◯ US $48

1200 meters from conference site, rooms with bathroom, breakfast

Hotel ROYAL∗∗∗ (deposit: US $100)  ◯ US $65 ◯ US $80  —

600 meters from conference site, rooms with bathroom, air cond., breakfast

I share a room ◯ with ◯ with anyone. I don’t want to share a room with anyone ◯.

CONDITIONS: cf. First Announcement

I ◯ enclose ◯ will send by May 15 ◯ a money order/certified cheque ◯ a copy of bank transfer of US $ , which includes US $ for registration fee and US $ for ho(s)tel

deposit.

Requests/remarks: